OpenStudy (anonymous):

Verify that cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

I changed the sec^4x

OpenStudy (anonymous):

i did (sec^2x)^2 which is (1+tan^2x)(1+tan^2x)

OpenStudy (anonymous):

and then 1+2tan^2x+tan^4x

OpenStudy (anonymous):

and then i got stuck

OpenStudy (anonymous):

lets also try right hand side

OpenStudy (anonymous):

ok what do you mean?

OpenStudy (anonymous):

\[\cot x+2\tan x+\tan 3x=???\]

OpenStudy (anonymous):

well the only thing you can do is change cot to cos/sin

OpenStudy (anonymous):

consider c as cos and s as sin tell me if its not clear \[\frac{ c }{ s }+\frac{ 2s }{ c }=\frac{ c^2+2s^2 }{ sc }=\frac{ 1-s^2+2s^2 }{ sc }=\frac{ 1-s^2 }{sc }\]

OpenStudy (anonymous):

tan x=sinx /cosx cot x=cos x/sin x so I added them up

OpenStudy (anonymous):

what about tan^3x

OpenStudy (anonymous):

sin 3x /cos 3x

OpenStudy (anonymous):

ok so then 1-s^2/sc = cos^2/sinxcosx=cosx/sinx=cot?

OpenStudy (anonymous):

wait its 1+s^2/sc

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

now i'm lost

OpenStudy (anonymous):

lets start over I also lost myself there

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

\[\cot x \sec 4x=\cot x(1+\tan 4x)\] \[\cot x+\cot x \tan 4x=\cot x +\cot x(\frac{ \sin 4x }{ \cos 4x })\]

OpenStudy (anonymous):

you are missing the 2tanx

OpenStudy (anonymous):

2tan^2x

OpenStudy (anonymous):

\[\frac{\sin 4x}{\cos 4x}=\frac{2\cos2x \sin 2x}{2\cos^22x-1}\]

OpenStudy (anonymous):

above I started with LHS left hand side so there is no tan initilly

OpenStudy (anonymous):

i messed up i give up sry

OpenStudy (anonymous):

ok.....thanks